Calculate the inverse z-transform of the signal below.
e ^ {- 2x}
Answer #1

Given 22 XCZ) = f(x) = 6 ax By applying the Tayloo -az XCZ) = (00) Jer) 222= 8W C-azok E K=0 -z 18W ē 11 formula ก | K=0 2 -